# Improving Precision in Kinematic Weak Lensing with MIRoRS: Model-Independent Restoration of Reflection Symmetries

### v1

@@ Abstract
We present a novel, model-independent technique for fitting the cross-component of weak lensing shear, $\gamma_{\times}$, along a line of sight by combining kinematic and photometric measurements of a single lensed galaxy. Rather than relying on parametric models, we fit for the shear parameter that best transforms the velocity field to restore its underlying symmetries, while also incorporating photometric data for the change in position angle due to shear. We first validate our technique with idealized mock data, exploring the method’s response to variations in shear, position angle, inclination, and noise. On this idealized mock data, our combined kinematic and photometric model demonstrates superior performance compared to traditional parametric or kinematic-only approaches. Subsequently, we apply our method to a dataset of 358 halos from the Illustris TNG simulations, achieving a notable reduction in the uncertainty of $\gamma_{\times}$ to 0.039, marking a substantial improvement over previous analysis of the dataset with a parametric model. Finally, we introduce an outlier rejection method based on Moran’s I-test for spatial autocorrelation. Identifying and filtering out halos with spatially correlated residuals reduces the overall uncertainty to 0.028. Our results underscore the efficacy of combining kinematic and photometric data for weak lensing studies, providing a more precise and targeted measurement of shear along an individual line of sight.
1 Introduction
@@
Weak gravitational lensing provides a robust method for probing the mass distribution of the universe (see Bartelmann & Schneider (2001) for a comprehensive review and Bartelmann & Maturi (2017) for a more contemporary treatise). As the light from an extended background source travels through the potential well of a foreground lens, it is distorted such that the image observed carries information about the underlying mass distribution of the lens. This distortion manifests in two primary effects: convergence and shear. Convergence, typically denoted $\kappa$, produces an isotropic scaling of the image, maintaining the shape while scaling the size. Shear, denoted $\gamma$, distorts the shape of the image, elongating along one axis and compressing in an orthogonal direction. Determining the intrinsic size of a source is typically difficult, thus the primary information available is in the ellipticity and orientation of the image.
@@
In the weak lensing regime, these effects are small, typically with ellipticity distortions less than 1% (McKay et al., 2001). Meanwhile, the intrinsically wide distribution of unlensed galaxy orientation and ellipticity, or shape noise, overwhelms the signal of a single source. In traditional weak lensing, shape noise is overcome by using a statistical ensemble of sources to search for this systematic distortion. While the uncertainty in the shape of any single source is large, it is presumed that there is no preferential orientation of an unlensed source ensemble. Averaging over a large number of sources, of order $\sim 10^{4}$ in cluster-galaxy lensing (Umetsu, 2020; Umetsu et al., 2014) and up to $\sim 10^{7}$ in cosmic shear studies (Troxel et al., 2018), is required to achieve a desirable signal-to-noise ratio.
@@
With no a priori knowledge of the shape and orientation of an unlensed galaxy, traditional weak lensing will always be plagued by shape noise. However, some source galaxies possess sufficient symmetries in their velocity field such that the light observed is not completely devoid of information on the unlensed condition of the galaxy. From the equivalence principle, it is known that lensing effects are independent of photon wavelength. Thus, the velocity field of a galaxy will also show the effects of shearing. Under idealized conditions, an unlensed disk galaxy has perpendicular kinematic axes and displays symmetry in its velocity field along its major axis and antisymmetry along its minor axis. These symmetries are broken under shear and it was first recognized by Blain (2002) that this deviation produces an observable that can be used to infer the shear parameter from a single galaxy.
@@
This work was extended by Morales (2006) and significant progress has since been made in the subfield of Kinematic Lensing (henceforth KL). The most common approach involves fitting a parametric model of a sheared velocity field to observations. This approach was most recently explored by Donet & Wittman (2023, hereafter DW23), where an uncertainty in $\gamma_{\times}\approx 0.08$ was found across a sample of 358 halos from Illustris TNG simulations (Nelson et al., 2019). This uncertainty represents an improvement over traditional weak lensing methods, but it remains larger than the typical available signal, underscoring the need for further refinement.
@@
These parametric modeling techniques are reliant upon the form of the velocity profile chosen. True rotation curves do not often match such an idealized model. While KL offers potential improvement over traditional weak lensing techniques on a per-source basis, an ensemble of galaxies is still required to overcome the variability in galaxy kinematic morphology, or kinematic shape noise (Gurri et al., 2020). However, de Burgh-Day et al. (2015, dBD15) proposed a model-independent approach, using only the symmetry properties of a galaxy. Later, it was shown by DiGiorgio et al. (2021, DiG21) that explicitly including the offset in the position angle between the kinematic and photometric axes, in conjunction with a kinematic model, reduces statistical uncertainty in the inferred shear parameter by a factor of 2–6. We seek to refine and extend these works.
@@
Our approach takes inspiration from dBD15, but has some fundamental differences. Starting from an observed velocity field, we transform the field back to the source plane through translation, rotation, and lensing shear. When transformed with true parameter values, the symmetry (antisymmetry) under reflections across the major (minor) axis will be restored. Performing these reflections and then transforming back to the detector frame gives two new velocity fields, one for each symmetry transformation. These fields are then combined into a single nonparametric model and compared to the original data to compute a residual and a likelihood to be optimized. We also incorporate some of the methods of DiG21. In our likelihood function, we include a term from the difference between the observed photometric and modeled kinematic axes.
@@
We first characterize our approach using idealized mock data to explore its response and limitations. After the initial testing, we applied our model to a more realistic dataset, comprised of 358 halos from the Illustris TNG simulations. This is the same dataset used by DW23, allowing for a direct comparison between our method and their parametric approach. Finally, we demonstrate an objective method for discarding outliers whose velocity fields do not follow the idealized assumptions. We utilized a Moran’s I-test for spatial autocorrelation to search for patterns in the residuals indicating unmodeled physical effects.
@@
The paper is structured as follows: in §2 we will discuss weak lensing and its effects on observables, §3 will outline our approach, and §4 will cover results and discussion.
2 Weak Lensing
@@
The theory of gravitational lensing in general, and weak lensing in particular, is well-established in the literature (Bartelmann & Schneider, 2001; Bartelmann & Maturi, 2017; Umetsu, 2020). Light from a background source, following a geodesic, will be deflected as it travels through the potential well of a massive body. This deflection will produce a distorted image, encoding information from the intervening mass distribution. For the purpose of this paper, we will limit our discussion to the weak lensing regime of small deflections. We will also assume that the characteristic size of the lensing mass distribution is small compared to the distances along the geodesic and adopt a thin-lens approximation of the lens.
@@
where the prefactor $\mu=((1-\kappa)^{2}-\lvert\gamma\rvert^{2})^{-1}$ gives the magnification, and the magnitude of total shear is given by $\lvert\gamma\rvert^{2}=\gamma_{+}^{2}+\gamma_{\times}^{2}$. Examining this tensor gives insight into the effect of each component. Convergence produces an isotropic focusing of the image while both components of shear produce anisotropic effects. The first component, $\gamma_{+}$, stretches and compresses the image along the coordinate axes. This produces an elliptical image of an otherwise circular source, but is indistinguishable from the inherent ellipticity of the unlensed object. The second component, $\gamma_{\times}$, shears the object along the $y=x$ axis. These components can be expressed in terms of the total magnitude of shear and angle $\alpha$ between the shear vector and coordinate axes,
$\gamma_{\times}$ $=\lvert{\gamma}\rvert\sin(2\alpha)$ (3)
$\gamma_{+}$ $=\lvert{\gamma}\rvert\cos(2\alpha).$ (4)
We can simplify this transformation by recognizing that $\kappa$, producing only isotropic effects, will not impact the symmetry properties of the galaxy. Neglecting $\kappa$ gives
$\mathcal{A}=\begin{pmatrix}1-\gamma_{+}&-\gamma_{\times}\\ -\gamma_{\times}&1+\gamma_{+}\end{pmatrix},$ (5)
$\mathcal{A}^{-1}=\mu\begin{pmatrix}1+\gamma_{+}&\gamma_{\times}\\ \gamma_{\times}&1-\gamma_{+}\end{pmatrix}.$ (6)
Figure 1: Velocity field of a non-sheared galaxy (top) and galaxy with $\lvert\gamma_{\times}\rvert=0.12$ shear applied (bottom) from a lens along the $y=x$ line from the source. The sheared galaxy displays a notable loss of symmetry in its kinematic field as well as an apparent rotation in its morphology.
2.1 Distortion of Observables
Figure 2: Pre and post-lensing kinematic and photometric axes corresponding to an elliptical isophote overlaid on the sheared velocity field. Prior to lensing, both the kinematic and photometric axes are aligned (gray). After lensing, the isophote will remain elliptical with orthogonal photometric axes (black). For small values of $\gamma_{+}$, the transformation of the isophote can be well-modeled as an ellipse rotated by an angle $\theta$ as shown in Equation 17. The kinematic axes (red), on the other hand, do not remain orthogonal. The discrepancy between the kinematic and photometric major and minor axes is given by Equation 20 and Equation 21 respectively.
At the heart of the KL method lies the fact that under a weak lensing shear transformation, a disk galaxy will lose its inherent symmetries. As can be seen in Figure 1, the non-sheared field exhibits a symmetry in its velocity field about the major axis and an antisymmetry across the minor axis. After applying the cross-component of shear ($\gamma_{\times}=0.12$ in this case), these symmetries are both lost. We use this loss of symmetry to fit for the unknown shear parameter directly by finding the transformation of the sheared field that restores the symmetry.
We can also take advantage of the fact that the kinematic and photometric axes are transformed differently, as seen in Figure 2. Our approach largely mirrors the work of DiG21, correcting for some inconsistencies that do not affect their ultimate findings. An isophote of the unlensed galaxy can be modeled as an ellipse in the source (x,y) coordinates,
@@ 2.1 Distortion of Observables
where q is the ratio between minor and major axes, $q=b/a$, in the source plane. We are interested in the image produced in the detector coordinates, $(x^{\prime},y^{\prime})$. Transforming from the source plane to the observed image in the detector plane through Equation 5, $\bm{x}=\mathcal{A}\bm{x}^{\prime}$, gives
$x$ $=x^{\prime}(1-\gamma_{+})-y^{\prime}\gamma_{\times}$ (8)
$y$ $=y^{\prime}(1+\gamma_{+})-x^{\prime}\gamma_{\times}.$ (9)
The isophote in the detector plane is then
$1=q^{2}(x^{\prime}(1-\gamma_{+})-y^{\prime}\gamma_{\times})^{2}+(y^{\prime}(1+\gamma_{+})-x^{\prime}\gamma_{\times})^{2}.$ (10)
Note, $q$ is the axis ratio in the source plane and is not directly observable. We will address the observed axis ratio later, with Equation 16. In the weak lensing regime, we expect the shear parameters to be small, allowing us to expand and drop any quadratic terms. Combining like terms yields
@@
Next, we will approximate the transformation of the isophote as a rotation rather than a shear. Under a counterclockwise rotation Equation 7 becomes
$1=q^{\prime 2}(x^{\prime}\cos\Delta\theta+y^{\prime}\sin\Delta\theta)^{2}+(y^{\prime}\cos\Delta\theta-x^{\prime}\sin\Delta\theta)^{2}.$ (12)
Here, the observed axis ratio, $q^{\prime}$, is the same in both the source and detector frames. We denote the photometric rotation angle as $\Delta\theta$ to emphasize that this rotation is the change in photometric position angle due to lensing. In the weak lensing regime we expect this rotation angle to be small and thus use small angle approximations to first order,
@@
we see that $\gamma_{+}\approx 0$ for the transformation to be modeled as a rotation. For our purposes, this is achieved by orienting the the galaxy axes such that the shear vector aligns with $\gamma_{\times}$. In a real observation, $\gamma_{+}$ will typically be small, but non-zero. We address this in §4.2.1.
We can also now revisit Equation 5 and Equation 6. Under the current approximations of weak lensing, and by defining our coordinate system such that $\alpha=45^{\circ}$ and thus $\lvert{\gamma}\rvert=\gamma_{\times}$, the transformation matrices are then
[…]
@@
indicates the observed axis ratio, $q^{\prime}=q$, the source axis ratio. Finally, the $x^{\prime}y^{\prime}$ term gives us
@@
We can then solve for $\gamma_{\times}$ in terms of the difference in angle,
@@
3 Methods
3.1 Overview

### v2

@@ Abstract
We present a novel, model-independent technique for fitting the cross-component of weak lensing shear, $\gamma_{\times}$, along a line of sight by combining kinematic and photometric measurements of a single lensed galaxy. Rather than relying on parametric models, we fit for the shear parameter that best transforms the velocity field to restore its underlying symmetries, while also incorporating photometric data for the change in position angle due to shear. We first validate our technique with idealized mock data, exploring the method’s response to variations in shear, position angle, inclination, and noise. On this idealized mock data, our combined kinematic and photometric model demonstrates superior performance compared to traditional parametric or kinematic-only approaches. We also explore the effects of asymmetric warps and show that rotation direction can impart a small bias on the fit of $\gamma_{\times}$. Subsequently, we apply our method to a dataset of 358 halos from the Illustris TNG simulations, achieving a notable reduction in the uncertainty of $\gamma_{\times}$ to 0.039, marking a substantial improvement over previous analysis of the dataset with a parametric model. Finally, we introduce an outlier rejection method based on Moran’s $I$ test for spatial autocorrelation. Identifying and filtering out halos with spatially correlated residuals reduces the overall uncertainty to 0.028. Our results underscore the efficacy of combining kinematic and photometric data for weak lensing studies, providing a more precise and targeted measurement of shear along an individual line of sight.
I Introduction
@@
Weak gravitational lensing provides a robust method for probing the mass distribution of the Universe (see Bartelmann and Schneider (2001) for a comprehensive review and Bartelmann and Maturi (2017) for a more contemporary treatise). As the light from an extended background source travels through the potential well of a foreground lens, it is distorted such that the image observed carries information about the underlying mass distribution of the lens. This distortion manifests in two primary effects: convergence and shear. Convergence, typically denoted $\kappa$, produces an isotropic scaling of the image, maintaining the shape while scaling the size. Shear, denoted $\gamma$, distorts the shape of the image, elongating along one axis and compressing in an orthogonal direction. This shear can be described by two components: $\gamma_{+}$ elongates and compresses along the coordinate axes, while the cross-component $\gamma_{\times}$ operates along the line $y=x$. Determining the intrinsic size of a source is typically difficult; thus, the primary information available is in the ellipticity and orientation of the image.
@@
In the weak lensing regime, these effects are small, typically with ellipticity distortions $1\%$ (McKay et al., 2001). Meanwhile, the intrinsically wide distribution of unlensed galaxy orientation and ellipticity, or shape noise, overwhelms the signal of a single source. In traditional weak lensing, shape noise is overcome by using a statistical ensemble of sources to search for this systematic distortion. While the uncertainty in the shape of any single source is large, it is presumed that there is no preferential orientation of an unlensed source ensemble. Averaging over a large number of sources, of the order of $\sim 10^{4}$ in cluster-galaxy lensing (Umetsu, 2020; Umetsu et al., 2014) and up to $\sim 10^{7}$ in cosmic-shear studies (Troxel et al., 2018), is required to achieve a desirable signal-to-noise ratio.
@@
With no a priori knowledge of the shape and orientation of an unlensed galaxy, traditional weak lensing will always be plagued by shape noise. However, some source galaxies possess sufficient symmetries in their velocity field such that the light observed is not completely devoid of information on the unlensed condition of the galaxy. From the equivalence principle, it is known that lensing effects are independent of photon wavelength. Thus, the velocity field of a galaxy will also show the effects of shearing. Under idealized conditions, an unlensed disk galaxy has perpendicular kinematic axes and displays symmetry in its velocity field along its major axis and antisymmetry along its minor axis. These symmetries are broken under shear, and it was first recognized by Blain (2002) that this deviation produces an observable that can be used to infer the shear parameter from a single galaxy.
This work was extended by Morales (2006), and significant progress has since been made in the subfield of Kinematic Lensing [^1] We adopt the term Kinematic Lensing broadly, rather than referring to the exact methods introduced in Huff et al. (2013). Conceptually similar approaches have gone by many different names including direct shear mapping (de Burgh-Day et al., 2015), precision weak lensing (Gurri et al., 2020), and velocity field lensing (Donet and Wittman, 2023). (henceforth KL). Following these early developments, KL methods have evolved in two primary directions shaped by their scientific goals. One avenue, pursued largely in the context of cosmic-shear surveys, aims to extract ensemble-average shear using minimal spectroscopic information (Huff et al., 2013; Xu et al., 2023; Pranjal et al., 2023; Huang et al., 2025). A second approach focuses on using spatially resolved velocity fields of individual galaxies to obtain precise per-source shear constraints, enabling detailed studies of localized shear (de Burgh-Day et al., 2015; Gurri et al., 2020; DiGiorgio et al., 2021; Donet and Wittman, 2023).
@@
In both applications, the most common approach involves fitting a parametric model of a sheared velocity field to observations. This approach was most recently explored by Donet and Wittman (2023, hereafter DW23), where an uncertainty in the cross-component of shear [^2] In parametric modeling of an individual galaxy, the coordinate system is most often chosen to exactly align the unlensed galaxy axes with the coordinates such that $\gamma_{+}$ acts along the galaxy axes, and $\gamma_{\times}$ acts at $45^{\circ}$. With the exception of mock data intentionally created with a nonzero position angle, we align our observed galaxies with the coordinate axes to best preserve this same relation. It should be noted, however, that alignment in the observed frame does not exactly correspond to alignment in the unlensed galaxy frame. of $\gamma_{\times}\approx 0.08$ was found across a sample of 358 halos from Illustris TNG simulations (Nelson et al., 2019). This uncertainty represents an improvement over traditional weak lensing methods, but it remains larger than the typical available signal, underscoring the need for further refinement.
These parametric modeling techniques are reliant upon the form of the velocity profile chosen. True velocity fields do not often match such an idealized model, and one of the primary findings of DW23 was that galaxies deviate significantly from intrinsic circularity. This has an immediate consequence for the inference of shear along the galaxy axes ($\gamma_{+}$ if the unlensed galaxy axes are aligned with the coordinate axes), which typically relies on an assumption of circularity when viewed face-on to determine the intrinsic axis ratio of the galaxy via the Tully–Fisher relation (Huff et al., 2013). In contrast, shear applied at $45^{\circ}$ to the galaxy axes ($\gamma_{\times}$ in galaxy-aligned coordinates) induces a symmetry-breaking distortion in the velocity field that is directly observable and does not depend on any assumptions about the unlensed axis ratio. While KL offers potential improvement over traditional weak lensing techniques on a per-source basis, an ensemble of galaxies may still be required to overcome the variability in galaxy kinematic morphology, or kinematic shape noise (Gurri et al., 2020, 2021).
@@
To increase precision without reliance on a parametric model, de Burgh-Day et al. (2015, dBD15) proposed a model-independent approach, using only the symmetry properties of a galaxy to constrain the cross-component of shear. This reduces the number of fitting parameters required from $\sim 11$ for a full parametric disk model to just five (centroid position and velocity, position angle, and $\gamma_{\times}$). Later, it was shown by DiGiorgio et al. (2021, DiG21) that explicitly including the offset in the position angle between the kinematic and photometric axes, in conjunction with a kinematic model, reduces statistical uncertainty in the inferred shear parameter by a factor of 2–6. We seek to refine and extend these works, focusing exclusively on the cross-component of shear and adopting a fully model-independent framework.
@@
Our method, MIRoRS (Model-Independent Restoration of Reflection Symmetries), takes inspiration from dBD15, but has some fundamental differences. Starting from an observed velocity field, we transform the field back to the source plane through translation, rotation, and lensing shear. When transformed with true parameter values, the symmetry (antisymmetry) under reflections across the major (minor) axis will be restored. Performing these reflections and then transforming back to the detector frame gives two new velocity fields, one for each symmetry transformation. These fields are then combined into a single nonparametric model and compared to the original data to compute a residual and a likelihood to be optimized. We also incorporate some of the methods of DiG21. In our likelihood function, we include a term from the difference between the observed photometric and modeled kinematic axes.
@@
We first characterize our approach using idealized mock data to explore its response and limitations. After the initial testing, we applied our model to a more realistic dataset, comprising 358 halos from the Illustris TNG simulations. This is the same dataset used by DW23, allowing for a direct comparison between our method and their parametric approach. Finally, we demonstrate an objective method for discarding outliers whose velocity fields do not follow the idealized assumptions. We utilized a Moran’s $I$ test for spatial autocorrelation to search for patterns in the residuals indicating unmodeled physical effects.
@@
The paper is structured as follows: in §II we will discuss weak lensing and its effects on observables; §III will outline our approach; and §IV will cover results and discussion.
II Weak Lensing
@@
The theory of gravitational lensing in general, and weak lensing in particular, is well-established in the literature (Bartelmann and Schneider, 2001; Bartelmann and Maturi, 2017; Umetsu, 2020). Light from a background source, following a geodesic, will be deflected as it travels through the potential well of a massive body. This deflection will produce a distorted image, encoding information from the intervening mass distribution. For the purpose of this paper, we will limit our discussion to the weak lensing regime of small deflections. We will also assume that the characteristic size of the lensing mass distribution is small compared to the distances along the geodesic and adopt a thin-lens approximation of the lens.
@@
where the prefactor $\mu=((1-\kappa)^{2}-\lvert\gamma\rvert^{2})^{-1}$ gives the magnification, and the magnitude of total shear is given by $\lvert\gamma\rvert^{2}=\gamma_{+}^{2}+\gamma_{\times}^{2}$. Examining this tensor gives insight into the effect of each component. Convergence produces an isotropic focusing of the image while both components of shear produce anisotropic effects. The first component, $\gamma_{+}$, stretches and compresses the image along the coordinate axes. This produces an elliptical image of an otherwise circular source, but is indistinguishable from the inherent ellipticity of the unlensed object. The second component, $\gamma_{\times}$, shears the object along the $y=x$ axis.
II.1 Distortion of Observables
At the heart of the MIRoRS method lies the fact that under a weak lensing shear transformation, a disk galaxy will lose its inherent symmetries. As can be seen in Figure 1, the nonsheared field exhibits a symmetry in its velocity field about the major axis and an antisymmetry across the minor axis. After applying the cross-component of shear ($\gamma_{\times}=0.12$ in this case), these symmetries are both lost. We use this loss of symmetry to fit for the unknown shear parameter directly by finding the transformation of the sheared field that restores the symmetry.
Figure 1: Velocity field of a nonsheared galaxy (top panel) and galaxy with $\gamma_{\times}=0.12$ shear applied (bottom panel) from a lens along the $y=x$ line from the source. The sheared galaxy displays a notable loss of symmetry in its kinematic field as well as an apparent rotation in its morphology.
Figure 2: Pre- and post-lensing kinematic and photometric axes corresponding to an elliptical isophote overlaid on the sheared velocity field. Prior to lensing, both the kinematic and photometric axes are aligned (gray). After lensing, the isophote will remain elliptical with orthogonal photometric axes (black). For small values of $\gamma_{+}$, the transformation of the isophote can be well modeled as an ellipse rotated by an angle $\theta$ as shown in Equation 8. The kinematic axes (red), on the other hand, do not remain orthogonal. The discrepancy between the kinematic and photometric major and minor axes is given by Equation 11 and Equation 12 respectively.
We can also take advantage of the fact that the kinematic and photometric axes are transformed differently, as can be seen in Figure 2. Following the approach of DiG21, an isophote of the unlensed galaxy can be modeled as an ellipse in the source ($x$, $y$) coordinates,
@@ II.1 Distortion of Observables
where q is the ratio between minor and major axes, $q=b/a$, in the source plane. We are interested in the image produced in the detector coordinates, $(x^{\prime},y^{\prime})$, which can be found by applying the shear transformation, $\bm{x}=\mathcal{A}\bm{x}^{\prime}$. In the weak lensing regime, we expect the shear parameters to be small, allowing us to expand and drop any quadratic terms yielding
@@
Next, we will approximate the transformation of the isophote as a rotation rather than a shear. Again, in the weak lensing regime, we expect this rotation angle to be small and thus use small angle approximations to first order,
@@
Here, the observed axis ratio, $q^{\prime}$, is the same in both the source and detector frames. We denote the photometric rotation angle as $\Delta\theta$ to emphasize that this rotation is the change in photometric position angle due to lensing.
@@
we see that $\gamma_{+}\approx 0$ for the transformation to be modeled as a rotation. For our purposes, this is achieved by orienting the the galaxy axes such that the shear vector aligns with $\gamma_{\times}$. In a real observation, $\gamma_{+}$ will typically be small, but nonzero. We address this in §IV.2.1.
@@
indicates the observed axis ratio, $q^{\prime}=q$, the source axis ratio. While derived in a different manner, this matches the ultimate findings of Eq. 18 in DiG21. Finally, the $x^{\prime}y^{\prime}$ term gives us
@@
Further, we can solve for $\gamma_{\times}$ in terms of the difference in angle,
@@
We can also now revisit Equation 1 and Equation 2. Under the current approximations of weak lensing, and by defining our coordinate system such that $\lvert{\gamma}\rvert=\gamma_{\times}$, the transformation matrices are then
[…]
III Methods
III.1 Overview